Project Cost $125 million in 1999, CPI 2018$/1999$=1.50, Exchange rate 6.71 ¥/$
What did it cost in 2018 ¥?
1
Expert's answer
2020-04-30T09:53:41-0400

The cost of a project in 2018 will be:

"Cost2018= Cost1999\times CPI=125 mln\times 1.5=187.5 mln."

With the exchange rate of 6.71, the cost of a project in yuans will be:

"187.5 mln\times 6.71 = 1 258 125 000 yuans."

In 2018 the project will cost 1 258 125 000 yuans.